OpenStudy (anonymous):

The book shop recieved a shipment of identification books. 1/3 of themwere bird books, 1/2 of the rest of the books were about plants, and 1/4 of the remaining books were mammal books. The 12 books that were left were insect books. How many books were in the shipment?

OpenStudy (anonymous):

im not sure how to figure it out sorry :/

OpenStudy (anonymous):

12*4=48 48*2=96 96*3=288 288 total books

OpenStudy (anonymous):

i got 48 books:)

OpenStudy (anonymous):

let matricked show you how:)

OpenStudy (anonymous):

(1-1/4+ 3/4 of the remainder books were on insects but no of insects books were 12 now leaving books on birds and plants the number is 12 * 4/3 = 16 but 16 represents( 1- (1/3) -(1-1/3)*1/2))= 1/3 of the total shipment thus the shipment had 16/(1/3) = 48 books

OpenStudy (anonymous):

thank you so much everybody :))